An automobile manufacturer claims that a new Hybrid model will average at least 60 miles per gallon of gas. A sample of 40 cars and and a 5% level of significance will be used for this analysis, The population standard deviation is assumed to be 7 miles per gallon. What is the Probability of a Type II error if the "true" mean is 57 miles per gallon? What is the probability of making a Type I error?
Answer #1

Ans:

left tailed test

P(type I error)=alpha=0.05

sample mean cut off value=60-1.645*(7/sqrt(40))=58.179

When true mean=57

z=(58.179-57)/(7/sqrt(40))

z=1.065

P(Type II error)=P(z>1.065)=0.1433